The organizers of a conference in the Houston Convention Center are evaluating the possibility of setting up a computer area where attendees can check their e-mail on computers provided by the organization. There will be one common queue for all computers and only one person uses a computer at a time. On average there are 10 attendee arrivals per hour and the average time a person spends on the computer is 15 minutes. To ensure that waiting times are not too long, the organizers want to ensure that the utilization of the computers doesn’t exceed 90 percent. At least how many computers do they need to have?

4
2
5
3
Answer #1

Activity time = 15 minutes

inter-arrival time = 60/10 = 6 mins

u = p/m*a

15/6/0.9=2.78

The correct answer is option D: 3